OpenStudy (nbarrera):

Help me Check MY answer+MEDAL! Which of the following numbers is prime? A. 153 B. 143 C. 163 D. 133 My answer is C just checkin if im right.

OpenStudy (anonymous):

you are correct

OpenStudy (welshfella):

153 is divisible by 3 143 is divisible by 11 133 is divisible by 7

OpenStudy (anonymous):

correct http://openstudy.com/study#/updates/52f15ceae4b0ec51a7e4cf9a

OpenStudy (michele_laino):

hint: 143= 11*13 153= 51*3

OpenStudy (nbarrera):

Thank you

OpenStudy (welshfella):

yw

OpenStudy (anonymous):

yw

OpenStudy (michele_laino):

thank you!

OpenStudy (nbarrera):

wish i can medal u all